OpenStudy (anonymous):

TRIGONOMETRY through what angle in degrees must the pendulum with 40cm. long be rotated to travel an arclength of 5cm.? TRIGONOMETRY through what angle in degrees must the pendulum with 40cm. long be rotated to travel an arclength of 5cm.? @Mathematics

OpenStudy (anonymous):

.125 radian= 7.16 degree (approx)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

the arc length is 5 cm, and the radius is 40 cm the definition of the radian measure of an angle is \[\frac{\text{ arc length }}{\text {radius }}\]

OpenStudy (anonymous):

in your case it is \[\frac{5}{40}=0.125\]

OpenStudy (anonymous):

but the question asks for the angle measured in degrees, so you have to convert from radian measure to degree measure,and you do that by multiplying by 180 and dividing by \[\pi\]

OpenStudy (anonymous):

hmmm, thanks you:) im not that bad in math im just confused in terms of problem solving

OpenStudy (anonymous):

therefore your answer is \[0.125\times \frac{180}{\pi}\] and then a calculator to give the answer in decimal form

OpenStudy (anonymous):

so, what is the unit of theta?

OpenStudy (anonymous):

0.125 is the angle measured in radians

OpenStudy (anonymous):

7.1619... is the angle measured in degrees, which is what you are asked for here is the conversion http://www.wolframalpha.com/input/?i=5%2F40+*180%2Fpi

OpenStudy (anonymous):

im right:)) i get the same answer, im making myself confused.. THANK YOU

OpenStudy (anonymous):

glad you got it